OpenStudy (anonymous):

Multiplying fractions is as easy as 1 2 3 But Can you Multiply With and without cancellation? Explain how or why not.

OpenStudy (anonymous):

Yes because you can still multiply the numerators and denominators like (2/3)(4/5)= (2*4)/(3*5) = 8/15

OpenStudy (anonymous):

2/3 * 4/5?

OpenStudy (anonymous):

yes, that notation would be another way of multiplying fractions if it didn't get cut off like that